OpenStudy (thatlilmathgirl12xd):

negative 1.1 over negative 3 as a decimal

OpenStudy (thatlilmathgirl12xd):

\[\frac{ -1.1 }{ -3 }\] as a decimal please.

OpenStudy (jtvatsim):

Are you allowed to use a calculator, or do you have to do it by hand?

OpenStudy (thatlilmathgirl12xd):

Oh this is just me asking. For my knowledge not for school.

OpenStudy (jtvatsim):

OK, I see. : ) Well, then if it's just for your knowledge... let me ask you a question: Would you be able to find this as a decimal if they were both positive? \[\frac{1.1}{3}\]

OpenStudy (thatlilmathgirl12xd):

0.36

OpenStudy (jtvatsim):

Sure, I'm assuming that you've cut off some of the decimals of course. :)

OpenStudy (thatlilmathgirl12xd):

i think.

OpenStudy (thatlilmathgirl12xd):

yes I did.

OpenStudy (jtvatsim):

OK, and one last question have you learned what to do if you have two negative numbers that multiply together? Like -1 x -4 = ?

OpenStudy (thatlilmathgirl12xd):

a positive.

OpenStudy (jtvatsim):

Perfect! You now have everything you need to answer your original question.

OpenStudy (jtvatsim):

Just like multiplication, when two negative numbers are divided, the result is positive.

OpenStudy (jtvatsim):

So, \[\frac{-1.1}{-3} = \frac{1.1}{3} = 0.366666...\]

OpenStudy (jtvatsim):

Sorry, that was a long way of answering your question. But I hope that makes sense and you can feel how it's the same as multiplication. : )

OpenStudy (thatlilmathgirl12xd):

Thank you

OpenStudy (jtvatsim):

You're welcome